# Scaling Foundation Models for Radar Scene Understanding

### v1

@@ 1 Introduction
Autonomous driving systems require robust perception capabilities that operate reliably across diverse environmental conditions. While cameras and LiDAR have driven recent advances, their performance degrades significantly in adverse weather conditions like rain, fog, and darkness. Radar sensors provide robust, all-weather perception through these adverse weather and lighting conditions [radar-bw-1, radar-bw-2, radar-bw-3]. Their ability to directly measure range and velocity makes them indispensable complements to vision-based sensing in autonomous driving. Yet despite these advantages, current machine learning approaches [radsegnet, pointillism, bootstrap-radar, radnet, radatron] remain fragmented and task-specific. Each downstream task, like object detection, semantic segmentation, and occupancy prediction, employs distinct input encodings, architectural designs, and training objectives [radar-fm-single-chip, racformer]. This fragmentation results in learned representations that are non-transferable across tasks and fail to generalize to diverse driving scenarios.
@@
The root of this limitation lies in how radar perception has traditionally been framed. Conventional Radar-ML pipelines [4d-scene-flow-radar, radar-segmentation-wsss, radar-road-segmentation, rodnet] focus on narrow, isolated objectives with categorical supervision (e.g., bounding boxes for detection or class labels for segmentation). Such supervision lacks the semantic richness needed to capture the complex spatial relationships and contextual cues critical for autonomous driving, such as which lane a vehicle occupies, whether a pedestrian is crossing, or how traffic infrastructure relates to the ego vehicle’s trajectory.
@@
Recent advances in vision-language models [clip-s4, mask-clip, vlm-for-segmentation-review] have demonstrated that language grounding offers a scalable and semantically rich supervision signal. Unlike categorical labels, natural language inherently encodes object types, spatial relations, and contextual information (e.g., “a stopped bus 20 m ahead”). Foundation models such as CLIP [clip] and BLIP [blip] have shown that aligning visual representations with textual descriptions enables transferable features that generalize across diverse tasks and unseen categories. Thus, language acts as a universal label space, providing high-level semantics that unify multiple perception objectives under a single representational framework [sam, sam-2].
@@
Translating this paradigm to radar presents both opportunities and challenges. On one hand, language supervision can guide radar encoders to emphasize semantically meaningful objects and their spatial relationships rather than raw signal patterns. By aligning radar embeddings with descriptive text, we impose semantic structure on radar features. Early explorations [bootstrap-radar, radar-clip-captioning] have provided preliminary evidence that radar-language alignment can improve downstream scene understanding. On the other hand, existing contrastive learning frameworks rely on binary matching between paired modalities, treating each element in the batch as either perfectly aligned or completely mismatched. This formulation is fundamentally flawed for radar scene understanding, where spatial configurations exist on a continuous spectrum of similarity. This prevents models from learning fine-grained spatial semantics. Training foundation models also requires large-scale, well-annotated datasets with rich textual descriptions, but real-world data collection at scale is quite expensive and time-consuming, posing a significant barrier.
@@
We introduce RadarFM, a radar foundation model that integrates language grounding to achieve semantically rich radar scene understanding. To address the data challenge, we leverage the CARLA simulator integrated with a realistic radar sensor [c-shenron-demo], enabling large-scale generation of diverse, well-annotated radar data across varied autonomous driving scenarios.
@@ 2 Related Works
Existing Radar Perception Approaches fall into three categories. Task-specific supervised methods train specialized architectures from scratch for individual perception tasks like object detection, segmentation, or scene flow [rodnet, radatron, radnet, k-radar, 4d-scene-flow-radar, racformer, radar-fm-single-chip]. Cross-modal supervised methods exploit camera and LiDAR supervision to generate pseudo-labels for radar training [radsegnet, radar-road-segmentation, 4d-scene-flow-radar]. Multi-sensor fusion architectures combine radar with cameras or LiDAR to leverage complementary sensor strengths [racformer, pointillism]. Unlike these fragmented approaches, we propose learning unified radar representations through structured language supervision that captures semantic richness without relying on task-specific architectures or cross-modal dependencies.
Vision-Language Foundation Models have demonstrated the power of natural language acting as a supervision signal for learning transferable visual representations. Contrastive pretraining methods align image and text embeddings in a shared space: CLIP [clip] achieves zero-shot transfer across 30+ vision tasks, while [align-google, coca, florence] scale to billions of noisy image-text pairs to learn robust cross-modal representations. Recent works [glip, region-clip, mask-clip, clip-s4] extend these models to dense prediction tasks through language-aware pretraining, and enable class-free semantic segmentation. These advances collectively demonstrate that language acts as a universal supervision signal [sam, sam-2, vlm-for-segmentation-review], enabling models to learn spatially-grounded semantic representations that generalize across tasks without task-specific annotations. Inspired by this paradigm, we propose to leverage language supervision for learning unified radar scene representations.
CLIP Loss Variants: The original CLIP [clip] framework employs a contrastive loss that treats image-text matching as a binary classification problem. While effective for generic image-level understanding, this binary formulation has inspired numerous extensions [soft-clip-loss, supervised-clip-loss, proto-nce] like relaxing the one-to-one constraint via soft targets, leveraging label information to pull same-class clusters together while pushing different classes apart, and cluster-level operations through learned prototypes. Some other variants [sigclip, slip, spatial-clip] replace softmax normalization with pairwise sigmoid loss, or combine self-supervised and language supervision in a multi-task framework. However, these variants focus on general-purpose vision tasks and fail to address the continuous spectrum of spatial similarity inherent to our problem setup. Our solution addresses this gap by quantifying scene similarity based on spatial configuration overlap, thus enabling fine-grained learning.
@@ 3 Dataset Curation
Training foundation models requires large-scale datasets with rich and accurate annotations. We leverage the CARLA simulator [carla-sim], a well-established platform for autonomous driving research, which provides perfect access to ground truth data for all actors in the scene. This simulator-based approach enables us to collect radar data with precise spatial annotations at scale. Since CARLA’s native radar sensor is highly simplified, we use an open-source implementation [c-shenron-demo] which models the propagation behavior of 77 GHz automotive radar very accurately.
@@ 4 RadarFM: Architecture and Approach
Our framework, RadarFM, consists of two key components: a CLIP-based [clip, vit] Vision encoder for radar heatmaps and a GPT-2-like Transformer [gpt2, transformers] encoder for captions, (2) hash-aware contrastive learning that accounts for continuous scene similarity. Additionally, we introduce a generative captioning capability that serves as a diagnostic tool to validate the learned representations.
@@ 4.1 Pre-training Architecture
We adopt the pretrained ViT-B/16 encoder from CLIP [clip], leveraging its robust visual feature extraction capabilities developed on large-scale image-text data. Given a radar range-angle heatmap, we replicate it across three channels to match CLIP’s expected input format. The heatmap is partitioned into $16\times 16$ patches with learnable positional embeddings, and the classification token output $\mathbf{z}_{r}\in\mathbb{R}^{512}$ serves as our radar scene representation (Figure 3).
@@ 4.2 Hash-Aware Contrastive Learning
Standard contrastive learning frameworks like CLIP [clip] treat sample pairs as binary: a matched image-text pair is positive (label = 1), while all other pairs in the batch are negative (label = 0). This binary formulation is suboptimal for radar scene understanding because scene similarity lies on a continuous spectrum (see the supplementary material for examples). Having binary labels will harshly penalize the model from learning fine-grained spatial distinctions, and instead drives it toward coarse, keyword-based matching.
@@ 4.3 Generative Captioning for Validation
Does contrastive pretraining truly learn spatially-grounded semantic understanding? To validate that RadarFM captures genuine scene comprehension, we extend our framework with generative captioning capabilities (Figure 4). Taking inspiration from ClipCap [clipcap], we train a lightweight projection mechanism from our pretrained radar encoder to GPT-2’s [gpt2] embedding space. This generative task serves as a diagnostic tool that assesses whether the model can produce accurate, detailed captions from radar inputs alone.
@@
Lightweight Mapping Network: The transformer-based mapping network $f_{\theta}$ projects the text embeddings into the input space of GPT-2 [gpt2]. This network generates expressive prefix embeddings through cross-attention mechanisms [clipcap], which is critical when the language model remains frozen, as the mapping network must encode all spatial information into a fixed-length prefix that can effectively steer GPT-2’s generation process. These prefix embeddings act as a soft prompt that conditions the auto-regressive generation. This design also substantially reduces the trainable parameters.
@@
where $c_{t}$ is the $t$-th token in the ground-truth caption, and $T$ is the caption length. This follows from the typical teacher-forcing [teacher-forcing] training methodology.
@@ 5.1 Pre-training Quality: Attention Analysis
To validate that hash-aware contrastive learning produces spatially-grounded representations, we analyze the attention patterns learned by the frozen ViT encoder. We employ attention rollout [attention-rollout], which recursively multiplies attention weights across transformer layers to reveal cumulative attention flow from the classification token to input patches.
@@ 5.3 Metrics
While standard captioning metrics such as BLEU [bleu], METEOR [meteor], and CIDEr [cider] effectively measure n-gram overlap, they are fundamentally flawed to assess the spatial accuracy. These metrics treat captions as bags of words or phrases, rewarding lexical similarity without verifying whether the spatial information is correct. To address this critical gap, we propose novel localization-aware metrics that directly quantify spatial reasoning accuracy by measuring how precisely the model predicts vehicle positions in the radar’s native coordinate system.
@@ 6 Conclusions
7 Supplementary Materials
This supplementary document provides additional details, analyses, and baseline comparisons to support the main paper.
8 Dataset Details and Statistics
8.1 CARLA Simulator Configuration and Data Collection
We leverage the open-source radar simulator implementation from [c-shenron-demo] to generate high-fidelity radar data across diverse driving scenarios. Our data collection spans all daylight conditions including morning, afternoon, evening, dusk, and night, following the scenario distribution from [neat]. In total, we collected 861,000 radar-caption frame pairs. The radar sensor is configured to simulate a 77 GHz Texas Instruments cascaded automotive radar system with 86 virtual antennas, providing realistic range-azimuth-Doppler measurements that accurately model real-world automotive radar characteristics.
8.2 Actor State Extraction and Filtering
For each simulation frame, we extract the complete state of all actors present in the CARLA world. Actors include vehicles, pedestrians (walkers), and traffic infrastructure elements such as traffic lights and signs. For each actor, we record: unique identifier, 3D location coordinates, transform (pitch, yaw, roll), velocity vector, acceleration vector, traffic light association status, current and ground truth traffic light states, traffic sign type, and applicable speed limits. This comprehensive state representation enables precise spatial reasoning and caption generation.
To focus on perception-relevant objects, we apply distance-based filtering relative to the ego vehicle’s position. Actors located beyond 40 meters from the ego vehicle are excluded from caption generation, as they fall outside the typical sensing and decision-making range for autonomous driving scenarios. For lane-wise classification, we transform actor positions into the ego vehicle’s local coordinate frame where the ego’s forward direction defines the positive x-axis and the left direction defines the positive y-axis. Each vehicle is classified based on its motion direction relative to the ego vehicle using dot products of forward vectors: vehicles are labeled as "same direction" (dot product >0.5>0.5), "opposing direction" (dot product <0.5<-0.5), or "other" (perpendicular or crossing traffic). Within the "same direction" category, lateral offsets determine lane assignment (in-lane, left-lane, or right-lane) using a configurable lane width threshold of 3 meters. Additionally, each vehicle is classified as being ahead, behind, or directly beside the ego vehicle based on the dot product between the ego’s forward vector and the relative position vector to that vehicle. This ego-centric spatial encoding provides the foundation for generating structured, spatially-grounded scene descriptions.
8.3 Natural Language Caption Generation
We generate natural language captions from the structured JSON representations using GPT-4.1 mini via the OpenAI API. Rather than relying on fixed templates, we prompt the language model to produce varied descriptions that maintain semantic consistency while introducing lexical diversity. The prompting structure consists of two components: a system prompt that instructs the model to adopt the perspective of a driver describing the scene to a passenger, and a user prompt that provides the JSON data along with detailed key-explanation mappings.
8.3.1 System Prompt
The system prompt establishes the generation style and constraints:
"You are an expert at describing traffic scenes from the perspective of a driver talking to a passenger. Only output the final caption. Do not include explanations. Also do not use short word forms with apostrophe, use the full expansion of words only."
This prompt ensures naturalistic descriptions while preventing the model from generating meta-commentary or using contractions, which could introduce inconsistencies during training.
8.3.2 Key-Explanation Mappings
The user prompt includes comprehensive mappings that provide semantic context for each spatial encoding element. These mappings are organized into two categories: distance bins and angular sectors.
Distance Bin Mappings:
  • 0-10m: "Objects within 0 to 10 meters from the ego vehicle. These are very close and may require immediate attention."
  • 10-20m: "Objects within 10 to 20 meters from the ego vehicle. These are close and should be monitored closely."
  • 20-30m: "Objects within 20 to 30 meters from the ego vehicle. These are at a moderate distance and should be noted."
  • 30-40m: "Objects within 30 to 40 meters from the ego vehicle. These are at a safe distance but still relevant."
Angular Sector and Feature Mappings:
  • total_vehicles: "Total number of vehicles in the distance bin."
  • left_lane_front_side: "Vehicles in the left adjacent lane ahead of the ego vehicle."
  • left_side: "Vehicles directly to the left side of the ego vehicle."
  • left_lane_back_side: "Vehicles in the left adjacent lane behind the ego vehicle."
  • in_lane_front_side: "Vehicles directly ahead of the ego vehicle in the same lane."
  • in_lane_back_side: "Vehicles directly behind the ego vehicle in the same lane."
  • right_lane_front_side: "Vehicles in the right adjacent lane ahead of the ego vehicle."
  • right_side: "Vehicles directly to the right side of the ego vehicle."
  • right_lane_back_side: "Vehicles in the right adjacent lane behind the ego vehicle."
  • opposing_lane_front: "Vehicles in the opposing lane ahead of the ego vehicle."
  • opposing_lane_back: "Vehicles in the opposing lane behind the ego vehicle."
  • other_lane_front: "Vehicles in perpendicular or intersecting lanes ahead of the ego vehicle."
  • other_lane_back: "Vehicles in perpendicular or intersecting lanes behind the ego vehicle."
  • applicable_traffic_signs: "Traffic signs that are applicable to the ego vehicle."
  • walkers: "Pedestrians in the scene."
8.3.3 Complete Prompt Example
To illustrate the complete prompting structure, consider a scene where the ego vehicle is on a highway with vehicles in multiple lanes. The structured JSON data is first preprocessed: numerical values are converted to word form (e.g., 3 becomes "three"), and keys within each distance bin are randomly shuffled to prevent the model from learning positional biases. The complete user prompt combines the JSON data with the key explanations:
Here is the JSON data:
  "0-10m": {
    "total_vehicles": "three",
    "in_lane_front_side": "one",
    "right_lane_back_side": "two"
  "10-20m": {
    "total_vehicles": "five",
    "right_lane_front_side": "three",
    "in_lane_back_side": "one",
    "right_lane_back_side": "one"
  "20-30m": {
    "total_vehicles": "four",
    "opposing_lane_front": "four"
  "30-40m": {
    "total_vehicles": "two",
    "in_lane_front_side": "one",
    "right_lane_front_side": "one"
  "applicable_traffic_signs": [],
  "walkers": "zero"
Here is the description of every key in the JSON data:
Distance Bin - 0-10m: Objects within 0 to 10 meters from the ego vehicle. These are very close and may require immediate attention.
[… additional distance bin descriptions …]
- total_vehicles: Total number of vehicles in the distance bin.
- in_lane_front_side: Vehicles directly ahead of the ego vehicle in the same lane.
- right_lane_back_side: Vehicles in the right adjacent lane behind the ego vehicle.
[… additional key descriptions …]
Given this data, generate natural language sentences describing the scene. Only provide the description. Do not explain how you arrived at it. Include all the fields provided even if is empty.
Additionally, we randomly vary whether to include the instruction "Do not add any distance bin numbers" to create captions with and without explicit distance references, further increasing linguistic diversity.
8.3.4 Caption Generation Parameters
To enhance caption diversity, we generate multiple caption variants for each radar frame using the following sampling parameters: temperature $T=0.5$ to balance creativity with consistency, top-p nucleus sampling with $p=0.9$ to constrain the probability mass while allowing varied word choices, maximum token length of 400 to accommodate detailed scene descriptions, and number of samples $n=5$ to provide multiple linguistic realizations of the same spatial structure. During training, we randomly sample one caption from this set for each radar frame, ensuring the model learns from diverse expressions of identical spatial configurations while maintaining semantic groundedness.
8.4 Hash Vector Encoding Specification
Each scene is encoded as a compact binary hash vector that enables efficient computation of spatial similarity during contrastive learning. The hash structure allocates bits hierarchically to capture both categorical information and spatial occupancy. Traffic sign states receive 4 bits encoding the presence of speed limits, traffic lights, stop signs, and yield signs as binary flags. Pedestrian counts are encoded using 3 bits, allowing representation of 0-7 walkers. The remaining bits encode vehicle presence across the spatial grid defined by distance bins and angular sectors.
For adjacent lane sectors (left-lane, in-lane, right-lane), we allocate 3 bits per sector per distance bin, encoding vehicle counts from 0 to 7. For opposing and perpendicular lane sectors, which typically contain fewer vehicles in structured driving scenarios, we allocate 5 bits per sector per distance bin, allowing counts up to 31. This differentiated bit allocation balances representational capacity with hash compactness. The complete hash vector for a scene with four distance bins totals: 4 (traffic signs) + 3 (walkers) + 12 sectors $\times$ (3 or 5 bits) $\times$ 4 bins = 55 bits for the 3-bit sectors and 40 bits for the 5-bit sectors, yielding a total of 103 bits per scene. Each distance bin’s hash value is constructed by bitwise OR operations that pack vehicle counts into the allocated bit positions, and the final hash list concatenates traffic sign bits, walker bits, and per-bin hash values in sequence.
9 Supplementary Videos
We provide supplementary video demonstrations that showcase RadarFM’s spatial understanding capabilities during continuous driving sequences. These videos present synchronized multi-modal views of the ego vehicle navigating through diverse traffic scenarios in the CARLA simulator. Each video frame displays: the camera reference view showing the visual scene from the driver’s perspective, the corresponding radar range-angle heatmap with attention rollout overlays highlighting spatially relevant regions, ground truth vehicle distributions in radial format across distance bins and angular sectors, and the model’s predicted caption describing the scene in natural language. The attention visualizations reveal how the learned representations focus on semantically meaningful spatial structures, while the frame-by-frame caption predictions demonstrate the model’s ability to track dynamic changes in traffic configurations as the ego vehicle progresses through intersections, lane changes, and multi-vehicle encounters. These videos provide qualitative evidence of RadarFM’s fine-grained spatial reasoning and its capacity to generate accurate, contextually appropriate scene descriptions across varying traffic densities and environmental conditions.

### v2

@@ 1 Introduction
Autonomous driving systems require robust perception capabilities that operate reliably across diverse environmental conditions. While cameras and LiDAR have driven recent advances, their performance degrades significantly in adverse weather conditions like rain, fog, and darkness. Radar sensors provide robust, all-weather perception through these adverse weather and lighting conditions [29, 28, 44]. Their ability to directly measure range and velocity makes them indispensable complements to vision-based sensing in autonomous driving. Yet despite these advantages, current machine learning approaches [1, 2, 12, 46, 24] remain fragmented and task-specific. Each downstream task, like object detection, semantic segmentation, and occupancy prediction, employs distinct input encodings, architectural designs, and training objectives [14, 5]. This fragmentation results in learned representations that are non-transferable across tasks and fail to generalize to diverse driving scenarios.
@@
The root of this limitation lies in how radar perception has traditionally been framed. Conventional Radar-ML pipelines [6, 16, 30, 40] focus on narrow, isolated objectives with categorical supervision (e.g., bounding boxes for detection or class labels for segmentation). Such supervision lacks the semantic richness needed to capture the complex spatial relationships and contextual cues critical for autonomous driving, such as which lane a vehicle occupies, whether a pedestrian is crossing, or how traffic infrastructure relates to the ego vehicle’s trajectory.
@@
Recent advances in vision-language models [13, 7, 10] have demonstrated that language grounding offers a scalable and semantically rich supervision signal. Unlike categorical labels, natural language inherently encodes object types, spatial relations, and contextual information (e.g., “a stopped bus 20 m ahead”). Foundation models such as CLIP [34] and BLIP [21] have shown that aligning visual representations with textual descriptions enables transferable features that generalize across diverse tasks and unseen categories. Thus, language acts as a universal label space, providing high-level semantics that unify multiple perception objectives under a single representational framework [18, 36].
@@
Translating this paradigm to radar presents both opportunities and challenges. On one hand, language supervision can guide radar encoders to emphasize semantically meaningful objects and their spatial relationships rather than raw signal patterns. By aligning radar embeddings with descriptive text, we impose semantic structure on radar features. Early explorations [12, 33] have provided preliminary evidence that radar-language alignment can improve downstream scene understanding. On the other hand, existing contrastive learning frameworks rely on binary matching between paired modalities, treating each element in the batch as either perfectly aligned or completely mismatched. This formulation is fundamentally flawed for radar scene understanding, where spatial configurations exist on a continuous spectrum of similarity. This prevents models from learning fine-grained spatial semantics. Training foundation models also requires large-scale, well-annotated datasets with rich textual descriptions, but real-world data collection at scale is quite expensive and time-consuming, posing a significant barrier.
@@
We introduce RadarFM, a radar foundation model that integrates language grounding to achieve semantically rich radar scene understanding. To address the data challenge, we leverage the CARLA simulator integrated with a realistic radar sensor [37, 25], enabling large-scale generation of diverse, well-annotated radar data across varied autonomous driving scenarios.
@@ 2 Related Works
Existing Radar Perception Approaches fall into three categories. Task-specific supervised methods train specialized architectures from scratch for individual perception tasks like object detection, segmentation, or scene flow [40, 24, 46, 31, 6, 5, 14]. Cross-modal supervised methods exploit camera and LiDAR supervision to generate pseudo-labels for radar training [1, 30, 6]. Multi-sensor fusion architectures combine radar with cameras or LiDAR to leverage complementary sensor strengths [5, 2]. Unlike these fragmented approaches, we propose learning unified radar representations through structured language supervision that captures semantic richness without relying on task-specific architectures or cross-modal dependencies.
Vision-Language Foundation Models have demonstrated the power of natural language acting as a supervision signal for learning transferable visual representations. Contrastive pretraining methods align image and text embeddings in a shared space: CLIP [34] achieves zero-shot transfer across 30+ vision tasks, while [15, 42, 43] scale to billions of noisy image-text pairs to learn robust cross-modal representations. Recent works [23, 47, 7, 13] extend these models to dense prediction tasks through language-aware pretraining, and enable class-free semantic segmentation. These advances collectively demonstrate that language acts as a universal supervision signal [18, 36, 10], enabling models to learn spatially-grounded semantic representations that generalize across tasks without task-specific annotations. Inspired by this paradigm, we propose to leverage language supervision for learning unified radar scene representations.
CLIP Loss Variants: The original CLIP [34] framework employs a contrastive loss that treats image-text matching as a binary classification problem. While effective for generic image-level understanding, this binary formulation has inspired numerous extensions [11, 17, 22] like relaxing the one-to-one constraint via soft targets, leveraging label information to pull same-class clusters together while pushing different classes apart, and cluster-level operations through learned prototypes. Some other variants [45, 27, 41] replace softmax normalization with pairwise sigmoid loss, or combine self-supervised and language supervision in a multi-task framework. However, these variants focus on general-purpose vision tasks and fail to address the continuous spectrum of spatial similarity inherent to our problem setup. Our solution addresses this gap by quantifying scene similarity based on spatial configuration overlap, thus enabling fine-grained learning.
@@ 3 Dataset Curation
Training foundation models requires large-scale datasets with rich and accurate annotations. We leverage the CARLA simulator [9], a well-established platform for autonomous driving research, which provides perfect access to ground truth data for all actors in the scene. This simulator-based approach enables us to collect radar data with precise spatial annotations at scale. Since CARLA’s native radar sensor is highly simplified, we use an open-source implementation [37, 25] which models the propagation behavior of 77 GHz automotive radar very accurately.
@@ 4 RadarFM: Architecture and Approach
Our framework, RadarFM, consists of two key components: a CLIP-based [34, 8] Vision encoder for radar heatmaps and a GPT-2-like Transformer [35, 38] encoder for captions, (2) hash-aware contrastive learning that accounts for continuous scene similarity. Additionally, we introduce a generative captioning capability that serves as a diagnostic tool to validate the learned representations.
@@ 4.1 Pre-training Architecture
We adopt the pretrained ViT-B/16 encoder from CLIP [34], leveraging its robust visual feature extraction capabilities developed on large-scale image-text data. Given a radar range-angle heatmap, we replicate it across three channels to match CLIP’s expected input format. The heatmap is partitioned into $16\times 16$ patches with learnable positional embeddings, and the classification token output $\mathbf{z}_{r}\in\mathbb{R}^{512}$ serves as our radar scene representation (Figure 3).
@@ 4.2 Hash-Aware Contrastive Learning
Standard contrastive learning frameworks like CLIP [34] treat sample pairs as binary: a matched image-text pair is positive (label = 1), while all other pairs in the batch are negative (label = 0). This binary formulation is suboptimal for radar scene understanding because scene similarity lies on a continuous spectrum (see the supplementary material for examples). Having binary labels will harshly penalize the model from learning fine-grained spatial distinctions, and instead drives it toward coarse, keyword-based matching.
@@ 4.3 Generative Captioning for Validation
Does contrastive pretraining truly learn spatially-grounded semantic understanding? To validate that RadarFM captures genuine scene comprehension, we extend our framework with generative captioning capabilities (Figure 4). Taking inspiration from ClipCap [26], we train a lightweight projection mechanism from our pretrained radar encoder to GPT-2’s [35] embedding space. This generative task serves as a diagnostic tool that assesses whether the model can produce accurate, detailed captions from radar inputs alone.
@@
Lightweight Mapping Network: The transformer-based mapping network $f_{\theta}$ projects the text embeddings into the input space of GPT-2 [35]. This network generates expressive prefix embeddings through cross-attention mechanisms [26], which is critical when the language model remains frozen, as the mapping network must encode all spatial information into a fixed-length prefix that can effectively steer GPT-2’s generation process. These prefix embeddings act as a soft prompt that conditions the auto-regressive generation. This design also substantially reduces the trainable parameters.
@@
where $c_{t}$ is the $t$-th token in the ground-truth caption, and $T$ is the caption length. This follows from the typical teacher-forcing [19] training methodology.
@@ 5.1 Pre-training Quality: Attention Analysis
To validate that hash-aware contrastive learning produces spatially-grounded representations, we analyze the attention patterns learned by the frozen ViT encoder. We employ attention rollout [3], which recursively multiplies attention weights across transformer layers to reveal cumulative attention flow from the classification token to input patches.
@@ 5.3 Metrics
While standard captioning metrics such as BLEU [32], METEOR [20], and CIDEr [39] effectively measure n-gram overlap, they are fundamentally flawed to assess the spatial accuracy. These metrics treat captions as bags of words or phrases, rewarding lexical similarity without verifying whether the spatial information is correct. To address this critical gap, we propose novel localization-aware metrics that directly quantify spatial reasoning accuracy by measuring how precisely the model predicts vehicle positions in the radar’s native coordinate system.